# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 5 LLM-Based autonomous agents in scientific research
To better promote new ideas regarding AI in scientific research, Jablonka et al. 447 organized a hackathon in March 2023. During this one-day hackathon, 14 innovative projects were developed to address chemical problems centered around predictive modeling, automation, knowledge extraction, and education. Below, we highlight agent-based approaches from that Hackathon. First, MAPI_LLM441 is an agent with access to the Materials Project API (MAPI) database that receives a query asking for a property of a material and then retrieves the relevant information from the dataset. If the material is not available on MAPI, the agent can search for similar materials and use in-context learning (ICL) to provide a prediction of the requested property. Additionally, MAPI_LLM also has a reaction module for synthesis proposal. Second, Rankovic et al. 439 used LLMs to make BO algorithms more accessible to a broader group of scientists; BOLLaMa implements a natural language interface to easily interact with BO software developed by their group448. Third, and similar to Ramos et al. 442 and Ranković and Schwaller 446 who employed LLMs in BO, Weiser et al. 449 focused on genetic algorithms (GA), a different optimization algorithm. In GA, pieces of information are stochastically combined and evaluated to guide the algorithm during the optimization. For chemistry, these pieces are often molecular fragments that are combined to compose a final whole molecular structure. Thus, Weiser et al. 449 used LLMs to implement common GA operators under the hypothesis that LLMs can generate new combined molecules better than random cross-over due to their sensory memory. Fourth, InsightGraph450 can draw general relationships between materials and their properties from JSON files. Circi and Badhwar 450 showed that LLMs can understand the structured data from a JSON format and reorganize the information in a knowledge graph. Further refinement of this tool could automate the process of describing relationships between materials across various scientific reports, a task that remains labor-intensive today. Fifth, Kruschwitz et al. 440 used ICL and LLMs to accurately predict the compressive strength of concrete formulations; Text2Concrete achieved predictive accuracy comparable with a Gaussian process regression (GPR) model, with the advantage that design principles can be easily added as context. This model was successfully applied in a BO algorithm following Ramos et al. 442 approach.428 For education purposes, multiple authors have raised the discussion about how LLMs can be used to support educators’ and instructors’ daily work.451, 452, 453, 454, 455 Finally, in this direction, Mouriño et al. 438 developed i-Digest, an agent whose perception module can understand audio tracks and video recordings. These audio recordings are transcribed to text using the Whisper456 model, and therefore, i-Digest is a digital tutor that generates questions to help students test their knowledge about the course material. These are just a few examples to showcase the capabilities of AI systems to innovate and generate solutions rapidly.
@@
More recently, Ma et al. 426 showed that agents can be trained to use tools. SciAgent426 was developed under the premise that finetuning LLMs for domain-specific applications is often impractical. Nevertheless, the agent can be fine-tuned with a set of tools that will enable them to perform well in a domain-specific task. These tools, typically Python functions, enable SciAgent to plan, retrieve, and use these tools to facilitate reasoning and answer domain-related questions effectively. The benchmark developed for SciAgent, known as SciToolBench, includes five distinct domains, each equipped with a set of questions and corresponding tools. The development of its retrieval and planning modules involved finetuning different LLMs on the MathFunc benchmark, resulting in a notable performance improvement of approximately $\sim$ 20% across all domains in SciToolBench compared to other LLMs.
@@ 5.1 Agents for literature review
PaperQA introduces a robust model that significantly reduces misinformation while improving the efficiency of information retrieval. This agent retrieves papers from online scientific databases, reasons about their content, and performs question-answering (QA) tasks. Its mechanism involves three primary components—“search”, “gather_evidence”, and “answer_question” and the authors adapted the Retrieval-Augmented Generation (RAG)457 algorithm to include inner loops on each step. For instance, PaperQA can perform multiple rounds of search and gather_evidence if, with reflection, not have enough evidence has been acquired to successfully answer_question. Further validating its capabilities, a new benchmark called LitQA demonstrates that PaperQA not only meets but exceeds human performance in solving complex, real-world scientific questions. By applying the RAG technique to full-text scientific papers, PaperQA sets a new standard in QA capabilities, achieving human-like performance in curated datasets without hallucination or selecting irrelevant citations.430 To further investigate PaperQA performance, the authors developed a new benchmark called LitQA, which designs tasks that mimic the complexity of real scientific inquiry. The benchmark comprises 50 multiple-choice questions sourced from biomedical papers published post-September 2021, ensuring they were not part of the LLM training data. In this setting, PaperQA demonstrated a precision rate of 87.9% and surpassed typical human accuracy, scoring 69.5% compared to 66.8%.430
@@
Building on top of PaperQA, WikiCrow exemplifies the practical application of AI in generating concise and relevant Wikipedia-style summaries. The authors show that while 16% of a human-created Wikipedia article comprises irrelevant statements, WikiCrow displays irrelevant information only 3% of the time. Their system also added 5% more correct citations when compared with original articles. Moreover, thanks to its foundation in the PaperQA framework430, WikiCrow achieves remarkable cost-efficiency. The authors estimate that WikiCrow can accomplish in a few days what would take humans approximately 60,000 hours, or about 6.8 years, thereby underscoring its ability to rapidly produce extensive scientific content. This efficiency exemplifies the reliability and transformative potential of AI in content creation 431.
@@
Following a different approach, the STORM model also addressed the problem of writing Wikipedia-like summaries, where the STORM acronym represents the Synthesis of Topic Outlines through Retrieval and Multi-perspective questions.427 This approach implements a two-step procedure. First, STORM retrieves multiple articles on a topic and uses an LLM to integrate various perspectives into a cohesive outline. Second, this outline is used to write each section of the Wikipedia-like summary individually. To create the outline, multiple articles discussing the topic of interest are retrieved by an “expert” LLM, which processes each one to create $N$ perspectives. Each perspective is then fed to a “writer” LLM, and a conversation is initiated between writer and expert. Finally, the $N$ conversations are used to design the final outline. The outline and the set of references, accessed by RAG, are given to the writer LLM. The writer LLM is prompted to use these inputs to generate each section of the article sequentially. Following this, all sections are merged and refined to eliminate redundancies and enhance coherence. Upon human evaluation, STORM is reported to be $\sim$ 25% more organized and present $\sim$ 10% better coverage when compared to a pure RAG approach. However, it was also less informative than human-written Wikipedia pages, and STORM presented a transfer of internet-borne biases, producing emotional articles, which is a major concern.
Ultimately, the idea of an “AI scientist” who can generate new, relevant research questions (RQ) has been pursued. Aligned with this vision, Liu et al. 434 developed CoQuest, partially automating the brainstorming for new RQs process. This system uses a human-computing interface (HCI) to allow the agent to create new RQs that can be further enhanced by human feedback. They developed two strategies for RQ generation: breadth-first, where the agent generates multiple RQs simultaneously following the original user’s prompt, and depth-first, where multiple RQs are created sequentially, building on the top of the previously generated RQ. For each RQ generation, the agent implements a ReAct369 framework with tools for literature discovery, hypothesis proposition, refinement, and evaluation. Upon evaluation of 20 HCI doctoral researchers by a post-interaction survey, the breadth-first approach was preferred by 60% of the evaluators. Interestingly, despite the evaluators’ report that the breadth-first approach gave them more control and resulted in more trustworthy RQs, the depth-first had better scores for novelty and surprise. This difference might be caused by the fact that the depth-first uses its own RQ to iterate. This process can introduce new keywords not considered by the users.
@@ 5.2 Agents for chemical innovation
Transitioning from literature synthesis to practical chemistry applications, we next explore how LLM-based agents have proven their capabilities to revolutionize routine chemical tasks towards an acceleration of molecular discovery and scientific research. Agents are flexible entities capable of developing prompt-specific workflows and executing a plan toward accomplishing a specific task. ChemCrow45 introduced a significant shift in how LLMs would be applied in chemistry, given that LLMs alone do not access information outside of their training data nor can they directly perform chemistry-related tasks.
@@
By augmenting LLMs with common chemical tools, computational or robotic, ChemCrow automates a broad spectrum of routine chemical tasks, demonstrating a significant leap in LLM applicability. Under human evaluation, ChemCrow consistently outperformed GPT-4, achieving an accuracy score of 9.24/10 compared to 4.79/10.45 The developers of ChemCrow have also considered the ethical implications and potential risks associated with its capabilities. ChemCrow’s high potential could be misused and exploited for malicious objectives, and therefore the authors have implemented safety checks and guidelines to prevent such misuse, or “dual usage”. Additionally, they acknowledge that ChemCrow, relying on an LLM, may not always provide completely accurate answers due to gaps in its chemical knowledge. As such, they recommend careful and responsible use of the tool, along with thorough scrutiny of its outputs. In summary, while ChemCrow presents a powerful new chemical assistant45, oversight of its use is required, and this agent’s access to tools has been deliberately limited to enhance security and avoid misuse.
@@
Similarly to ChemCrow45, Chemist-X458 uses RAG to get up-to-date literature information and use it to reliably solve user’s questions. Nevertheless, Chemist-X focuses on designing chemical reactions to achieve a given molecule. It works in three phases: (1) First, the agent searches molecule databases for similar molecules, then (2) there is a look-up for online literature searching for chemical reactions capable of converting the list of similar molecules in the target. Lastly, (3) machine learning models are used to propose the reaction conditions. To validate their agent, the authors used Chemist-X to design an HTS experiment aiming to produce 6-(1-methyl-1H-indazol-4-yl), resulting in a maximum yield of 98.6%.
@@
On the other hand, the Coscientist46 system exemplifies the integration of semi-autonomous robots in planning, conceiving, and performing chemical reactions with minimal human intervention. At its core, the system features a main module named ‘PLANNER’, which is supported by four submodules. These submodules, or tools, are responsible for performing actions such as searching the web for organic synthesis, executing Python code, searching the hardware documentation, and performing a reaction in an automated lab.46 Utilizing this framework, the Coscientist successfully conducted two types of chemical coupling reactions, Suzuki-Miyaura and Sonogashira, in a semi-automated fashion, with manual handling of initial reagents and solvents. Additionally, Coscientist was also used to optimize reaction conditions. In contrast to Ramos et al. 442, who used LLMs within a Bayesian Optimization (BO) algorithm as a surrogate model, Boiko et al. 46 approached the optimization task as a strategic “game” aimed at maximizing reaction yield by selecting optimal reaction conditions. This demonstrates the ability of GPT-4 to effectively reason about popular chemical reactions – possibly via comprehensive coverage in pretraining. The authors have indicated that the code for their agent will be released following changes in U.S. regulations on AI and its scientific applications. At the time of writing, the code remains unreleased, but a simple example that calculates the square roots of random numbers has been provided to illustrate their approach.46. These examples underscore the transformative role of LLMs in enhancing and automating chemical processes, which will likely accelerate chemical discovery.
@@
Automated workflows in protein research have also been explored. ProtAgent429 is a multi-agent system designed to automate and optimize protein design with minimal human intervention. This system comprises three primary agents: Planner, Assistant, and Critic. The Planner is tasked with devising a strategy to address the given problem, the Assistant executes the plan using specialized tools and API calls, and the Critic supervises the entire process, providing feedback and analyzing outcomes. These agents collaborate through a dynamic group chat managed by a fourth agent, the Chat Manager. Tasks executed by this team include protein retrieval and analysis, de novo protein design, and conditioned protein design using Chroma459 and OmegaFold460.
@@
Similarly to ProtAgent, Liu et al. 425 created a team of AI-made scientists (TAIS) to scientific discovery without human intervention. However, their agents have roles analog to human roles in scientific discovery, for instance, project manager, data engineer, code reviewer, statistician, and domain expert. While in ProtAgent429 agents interact through the Chat Manager only, TAIS425 enables AI scientists to interact between themselves directly using pre-defined collaboration pipelines. To evaluate TAIS, the authors curated the Genetic Question Exploration (GenQEX) benchmark, which consists of 457 selected genetic data questions. As a case study, the authors show TAIS’s answer to the prompt ‘‘What genes are associated with Pancreatic Cancer when considering conditions related to Vitamin D Levels?’’. The system identified 20+ genes with a prediction accuracy of 80%.
@@
Innovation can also be achieved by looking into data from a different point-of-view to get new insights. Automating querying databases was investigated by Ramos et al. 441 with a ReAct agent with access to the MAPI dataset. This concept was extended by Chiang et al. 423. LLaMP423 is a RAG-based ReAct agent that can interact with MAPI, arXiv, Wikipedia, and has access to atomistic simulation tools. The authors showed that grounding the responses on high-fidelity information (a well-known dataset) enabled the agent to perform inferences without fine-tuning.
@@
The agents in chemistry, as exemplified by ChemCrow45 and Coscientist46, highlight a significant shift towards automation and enhanced efficiency in molecular discovery and scientific research. These systems demonstrate the potential of integrating LLMs with chemical tools and automation frameworks, achieving impressive accuracy and effectiveness in tasks ranging from routine chemical operations to complex reaction optimizations. Similarly, ProtAgent429 and TAIS425 systems showcase the versatility of multi-agent frameworks in automating protein design and genetic research, pushing the boundaries of what AI-driven scientific discovery can achieve. These studies collectively showcase the incredible potential of agents in chemical and biological research, promising automation of routine tasks, easing the application of advanced techniques and analyses, and accelerating discoveries. However, they also underscore the necessity for meticulous oversight and responsible development to harness their full potential while mitigating risks.461

### v2

@@ 5 LLM-Based autonomous agents in scientific research
To better promote new ideas regarding AI in scientific research, Jablonka et al. 473 organized a hackathon in March 2023. During this one-day hackathon, 14 innovative projects were developed to address chemical problems centered around predictive modeling, automation, knowledge extraction, and education. Below, we highlight agent-based approaches from that Hackathon. First, MAPI_LLM467 is an agent with access to the Materials Project API (MAPI) database that receives a query asking for a property of a material and then retrieves the relevant information from the dataset. If the material is not available on MAPI, the agent can search for similar materials and use in-context learning (ICL) to provide a prediction of the requested property. Additionally, MAPI_LLM also has a reaction module for synthesis proposal. Second, Rankovic et al. 465 used LLMs to make BO algorithms more accessible to a broader group of scientists; BOLLaMa implements a natural language interface to easily interact with BO software developed by their group474. Third, and similar to Ramos et al. 468 and Ranković and Schwaller 472 who employed LLMs in BO, Weiser et al. 475 focused on genetic algorithms (GA), a different optimization algorithm. In GA, pieces of information are stochastically combined and evaluated to guide the algorithm during the optimization. For chemistry, these pieces are often molecular fragments that are combined to compose a final whole molecular structure. Thus, Weiser et al. 475 used LLMs to implement common GA operators under the hypothesis that LLMs can generate new combined molecules better than random cross-over due to their sensory memory. Fourth, InsightGraph476 can draw general relationships between materials and their properties from JSON files. Circi and Badhwar 476 showed that LLMs can understand the structured data from a JSON format and reorganize the information in a knowledge graph. Further refinement of this tool could automate the process of describing relationships between materials across various scientific reports, a task that remains labor-intensive today. Fifth, Kruschwitz et al. 466 used ICL and LLMs to accurately predict the compressive strength of concrete formulations; Text2Concrete achieved predictive accuracy comparable with a Gaussian process regression (GPR) model, with the advantage that design principles can be easily added as context. This model was successfully applied in a BO algorithm following Ramos et al. 468 approach.454 For education purposes, multiple authors have raised the discussion about how LLMs can be used to support educators’ and instructors’ daily work.477, 478, 479, 480, 481 Finally, in this direction, Mouriño et al. 464 developed i-Digest, an agent whose perception module can understand audio tracks and video recordings. These audio recordings are transcribed to text using the Whisper482 model, and therefore, i-Digest is a digital tutor that generates questions to help students test their knowledge about the course material. These are just a few examples to showcase the capabilities of AI systems to innovate and generate solutions rapidly.
@@
More recently, Ma et al. 452 showed that agents can be trained to use tools. SciAgent452 was developed under the premise that finetuning LLMs for domain-specific applications is often impractical. Nevertheless, the agent can be fine-tuned with a set of tools that will enable them to perform well in a domain-specific task. These tools, typically Python functions, enable SciAgent to plan, retrieve, and use these tools to facilitate reasoning and answer domain-related questions effectively. The benchmark developed for SciAgent, known as SciToolBench, includes five distinct domains, each equipped with a set of questions and corresponding tools. The development of its retrieval and planning modules involved finetuning different LLMs on the MathFunc benchmark, resulting in a notable performance improvement of approximately $\sim$ 20% across all domains in SciToolBench compared to other LLMs.
@@ 5.1 Agents for literature review
PaperQA introduces a robust model that significantly reduces misinformation while improving the efficiency of information retrieval. This agent retrieves papers from online scientific databases, reasons about their content, and performs question-answering (QA) tasks. Its mechanism involves three primary components—“search”, “gather_evidence”, and “answer_question” and the authors adapted the Retrieval-Augmented Generation (RAG)483 algorithm to include inner loops on each step. For instance, PaperQA can perform multiple rounds of search and gather_evidence if, with reflection, not have enough evidence has been acquired to successfully answer_question. Further validating its capabilities, a new benchmark called LitQA demonstrates that PaperQA not only meets but exceeds human performance in solving complex, real-world scientific questions. By applying the RAG technique to full-text scientific papers, PaperQA sets a new standard in QA capabilities, achieving human-like performance in curated datasets without hallucination or selecting irrelevant citations.456 To further investigate PaperQA performance, the authors developed a new benchmark called LitQA, which designs tasks that mimic the complexity of real scientific inquiry. The benchmark comprises 50 multiple-choice questions sourced from biomedical papers published post-September 2021, ensuring they were not part of the LLM training data. In this setting, PaperQA demonstrated a precision rate of 87.9% and surpassed typical human accuracy, scoring 69.5% compared to 66.8%.456
@@
Building on top of PaperQA, WikiCrow exemplifies the practical application of AI in generating concise and relevant Wikipedia-style summaries. The authors show that while 16% of a human-created Wikipedia article comprises irrelevant statements, WikiCrow displays irrelevant information only 3% of the time. Their system also added 5% more correct citations when compared with original articles. Moreover, thanks to its foundation in the PaperQA framework456, WikiCrow achieves remarkable cost-efficiency. The authors estimate that WikiCrow can accomplish in a few days what would take humans approximately 60,000 hours, or about 6.8 years, thereby underscoring its ability to rapidly produce extensive scientific content. This efficiency exemplifies the reliability and transformative potential of AI in content creation 457.
@@
Following a different approach, the STORM model also addressed the problem of writing Wikipedia-like summaries, where the STORM acronym represents the Synthesis of Topic Outlines through Retrieval and Multi-perspective questions.453 This approach implements a two-step procedure. First, STORM retrieves multiple articles on a topic and uses an LLM to integrate various perspectives into a cohesive outline. Second, this outline is used to write each section of the Wikipedia-like summary individually. To create the outline, multiple articles discussing the topic of interest are retrieved by an “expert” LLM, which processes each one to create $N$ perspectives. Each perspective is then fed to a “writer” LLM, and a conversation is initiated between writer and expert. Finally, the $N$ conversations are used to design the final outline. The outline and the set of references, accessed by RAG, are given to the writer LLM. The writer LLM is prompted to use these inputs to generate each section of the article sequentially. Following this, all sections are merged and refined to eliminate redundancies and enhance coherence. Upon human evaluation, STORM is reported to be $\sim$ 25% more organized and present $\sim$ 10% better coverage when compared to a pure RAG approach. However, it was also less informative than human-written Wikipedia pages, and STORM presented a transfer of internet-borne biases, producing emotional articles, which is a major concern.
@@ 5.2 Agents for chemical innovation
Transitioning from literature synthesis to practical chemistry applications, we next explore how LLM-based agents have proven their capabilities to revolutionize routine chemical tasks towards an acceleration of molecular discovery and scientific research. Agents are flexible entities capable of developing prompt-specific workflows and executing a plan toward accomplishing a specific task. ChemCrow47 introduced a significant shift in how LLMs would be applied in chemistry, given that LLMs alone do not access information outside of their training data nor can they directly perform chemistry-related tasks.
@@
By augmenting LLMs with common chemical tools, computational or robotic, ChemCrow automates a broad spectrum of routine chemical tasks, demonstrating a significant leap in LLM applicability. Under human evaluation, ChemCrow consistently outperformed GPT-4, achieving an accuracy score of 9.24/10 compared to 4.79/10.47 The developers of ChemCrow have also considered the ethical implications and potential risks associated with its capabilities. ChemCrow’s high potential could be misused and exploited for malicious objectives, and therefore the authors have implemented safety checks and guidelines to prevent such misuse, or “dual usage”. Additionally, they acknowledge that ChemCrow, relying on an LLM, may not always provide completely accurate answers due to gaps in its chemical knowledge. As such, they recommend careful and responsible use of the tool, along with thorough scrutiny of its outputs. In summary, while ChemCrow presents a powerful new chemical assistant47, oversight of its use is required, and this agent’s access to tools has been deliberately limited to enhance security and avoid misuse.
@@
Similarly to ChemCrow47, Chemist-X484 uses RAG to get up-to-date literature information and use it to reliably solve user’s questions. Nevertheless, Chemist-X focuses on designing chemical reactions to achieve a given molecule. It works in three phases: (1) First, the agent searches molecule databases for similar molecules, then (2) there is a look-up for online literature searching for chemical reactions capable of converting the list of similar molecules in the target. Lastly, (3) machine learning models are used to propose the reaction conditions. To validate their agent, the authors used Chemist-X to design an HTS experiment aiming to produce 6-(1-methyl-1H-indazol-4-yl), resulting in a maximum yield of 98.6%.
@@
On the other hand, the Coscientist48 system exemplifies the integration of semi-autonomous robots in planning, conceiving, and performing chemical reactions with minimal human intervention. At its core, the system features a main module named ‘PLANNER’, which is supported by four submodules. These submodules, or tools, are responsible for performing actions such as searching the web for organic synthesis, executing Python code, searching the hardware documentation, and performing a reaction in an automated lab.48 Utilizing this framework, the Coscientist successfully conducted two types of chemical coupling reactions, Suzuki-Miyaura and Sonogashira, in a semi-automated fashion, with manual handling of initial reagents and solvents. Additionally, Coscientist was also used to optimize reaction conditions. In contrast to Ramos et al. 468, who used LLMs within a Bayesian Optimization (BO) algorithm as a surrogate model, Boiko et al. 48 approached the optimization task as a strategic “game” aimed at maximizing reaction yield by selecting optimal reaction conditions. This demonstrates the ability of GPT-4 to effectively reason about popular chemical reactions – possibly via comprehensive coverage in pretraining. The authors have indicated that the code for their agent will be released following changes in U.S. regulations on AI and its scientific applications. At the time of writing, the code remains unreleased, but a simple example that calculates the square roots of random numbers has been provided to illustrate their approach.48. These examples underscore the transformative role of LLMs in enhancing and automating chemical processes, which will likely accelerate chemical discovery.
@@
Automated workflows in protein research have also been explored. ProtAgent455 is a multi-agent system designed to automate and optimize protein design with minimal human intervention. This system comprises three primary agents: Planner, Assistant, and Critic. The Planner is tasked with devising a strategy to address the given problem, the Assistant executes the plan using specialized tools and API calls, and the Critic supervises the entire process, providing feedback and analyzing outcomes. These agents collaborate through a dynamic group chat managed by a fourth agent, the Chat Manager. Tasks executed by this team include protein retrieval and analysis, de novo protein design, and conditioned protein design using Chroma485 and OmegaFold486.
@@
Similarly to ProtAgent, Liu et al. 450 created a team of AI-made scientists (TAIS) to scientific discovery without human intervention. However, their agents have roles analog to human roles in scientific discovery, for instance, project manager, data engineer, code reviewer, statistician, and domain expert. While in ProtAgent455 agents interact through the Chat Manager only, TAIS450 enables AI scientists to interact between themselves directly using pre-defined collaboration pipelines. To evaluate TAIS, the authors curated the Genetic Question Exploration (GenQEX) benchmark, which consists of 457 selected genetic data questions. As a case study, the authors show TAIS’s answer to the prompt ‘‘What genes are associated with Pancreatic Cancer when considering conditions related to Vitamin D Levels?’’. The system identified 20+ genes with a prediction accuracy of 80%.
@@
Innovation can also be achieved by looking into data from a different point-of-view to get new insights. Automating querying databases was investigated by Ramos et al. 467 with a ReAct agent with access to the MAPI dataset. This concept was extended by Chiang et al. 447. LLaMP447 is a RAG-based ReAct agent that can interact with MAPI, arXiv, Wikipedia, and has access to atomistic simulation tools. The authors showed that grounding the responses on high-fidelity information (a well-known dataset) enabled the agent to perform inferences without fine-tuning.
@@
The agents in chemistry, as exemplified by ChemCrow47 and Coscientist48, highlight a significant shift towards automation and enhanced efficiency in molecular discovery and scientific research. These systems demonstrate the potential of integrating LLMs with chemical tools and automation frameworks, achieving impressive accuracy and effectiveness in tasks ranging from routine chemical operations to complex reaction optimizations. Similarly, ProtAgent455 and TAIS450 systems showcase the versatility of multi-agent frameworks in automating protein design and genetic research, pushing the boundaries of what AI-driven scientific discovery can achieve. These studies collectively showcase the incredible potential of agents in chemical and biological research, promising automation of routine tasks, easing the application of advanced techniques and analyses, and accelerating discoveries. However, they also underscore the necessity for meticulous oversight and responsible development to harness their full potential while mitigating risks.487